OpenStudy (anonymous):

Y= number of vats required. Write an equation using y that can be used to determine the total amount of an order. All orders are billed $30 for shipping, and all orders have a 6% sales tax added to the total order. Vats are $400 each. Can anyone walk me through this?

OpenStudy (anonymous):

$400y +6% + 30 = x ? is this it?

OpenStudy (amistre64):

y = (.06)x +30

OpenStudy (amistre64):

that might be wrong

OpenStudy (anonymous):

can u explain how u got there?

OpenStudy (amistre64):

just trying to parse the information... we know that 30 is added outraight y = ... +30

OpenStudy (amistre64):

one vat is 400 and the tax is 6% 24 dollars per vat..right?

OpenStudy (anonymous):

right...

OpenStudy (amistre64):

400 + 24 each vat 424x where x is the number of vats.... make sense so far?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

y=424x+30 but I wonder what "total" order means...if it includes that $30....probably not....

OpenStudy (anonymous):

the total amount of vats ordered

OpenStudy (amistre64):

id go with this then: y = 424x + 30

OpenStudy (anonymous):

by finding that number, i should be able to find the total dollar amount of the order

OpenStudy (amistre64):

yes, if they order 5 vats; then y = 424(5) + 30 = something......